OpenStudy (anonymous):

write function in terms of its cofunction... sin (alpha + 20 degrees) = cos...

OpenStudy (kc_kennylau):

\[\begin{array}{rl}\sin(90-\theta)&=&\cos\theta\\\sin\theta&=&\cos(90-\theta)\end{array}\]

OpenStudy (kc_kennylau):

You can use either formula, but I would recommend the second